# Optimization under rare events: scaling laws for linear chance-constrained programs

### v1

@@ 4. Heavy tails
Throughout this section, we suppose that $L$ has a distribution that is of multivariate regular variation, which covers the case in which the elements of $L$ have a power law. To describe our assumptions in a convenient form, we use a representation which (apart from not taking logarithms) is different from the one in Section 3, building on an equivalent characterization of regular variation developed in Basrak et al. (2002)). Specifically, let ||$\cdot$|| ⋅ | be the ${\mathcal{L}}_{1}$ norm. Then $\bar{F}(r)={\bf P}(|L|>r)=h(r)r^{-\alpha},\alpha>1$, with $h$ slowly varying (i.e. $h(ar)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for $a>0$) and
@@
The idea behind this representation is that, if we transform $L$ in polar coordinates, the radius and angle become independent, conditional upon the radius getting large. This formulation allows dependence between extreme values of risk factors. For background on multivariate regular variation, see Basrak et al. (2002), Resnick (2007).
We now construct an appropriately normalized version of ($\rm{CCP}_{\delta}$) in the regime $\delta\downarrow 0$. Set $x=y/r_{\delta}$, with $r_{\delta}:=\bar{F}^{-1}(\delta)=\inf\{r:{\bf P}(|L|>r)\leq\delta\}$. Now, write the constraint in ($\rm{CCP}_{\delta}$) as ${\bf P}(\phi(y,L)>r_{\delta})\leq\delta.$ We assume that $\bar{F}(t)$ is eventually strictly monotone and continuous which makes the constraint equivalent to
@@
maximize $\mbox{maximize }y^{T}c\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq 1.$ ($\rm{HT}$)
@@
[…]
Unlike in the light-tail case, the value $y^{*}$ leads to an asymptotically optimal solution as asymptotic feasibility is guaranteed. In contrast, in the heavy-tailed case ($\rm{CVaR}_{\delta}$) is no longer asymptotically optimal:
@@
$v_{{\bf CVaR},\delta}/v_{\delta}\rightarrow 1-1/\alpha$ as $\delta\downarrow 0$.
The reason behind the subobtimality is that the overshoot $\phi(y,L)-r_{\delta}$, conditioned on $\phi(y,L)>r_{\delta}$ is not of smaller order than $r_{\delta}$ when $\phi(y,L)$ has a heavy tail.
In the previous section, we observed that the sample approximation ($\mathrm{SCP}_{k}$) of ($\rm{CCP}_{\delta}$) has some optimality properties in the light-tail case, since the space of feasible solutions converges, after appropriate normalization, to a deterministic set which coincides with the feasible region of the limiting optimization problem ($\rm{LT}$). In the heavy-tail case, this is no longer true. In contrast, a random solution emerges, making any optimality property elusive.
To see this, consider the simplest possible case, where the goal is to maximize a scalar $x$ subject to the constraint that ${\bf}P(xL>1)\leq\delta$, with $L$ also scalar. ($\mathrm{SCP}_{k}$) now can be formulated as maximizing $x$ subject to $x\max_{i=1}^{k}L^{(i)}\leq 1$. If $L$ is light-tailed, we can normalize $\max_{i=1}^{k}L^{(i)}$ such that it converges to a deterministic constant (we have done this in a more general setting to obtain Theorem 3). However, in the heavy-tail case, such a normalization is not possible. In fact, $\max_{i=1}^{k}L^{(i)}/\bar{F}^{-1}(1/k)$ converges in distribution to a random variable with a Fréchet distribution, see Resnick (2007), Nair et al. (2022) for background.
To describe our general technical result we use point-process machinery commonly used in extreme-value theory, cf. Resnick (2007) as well as random set theory as developed in (Molchanov and Strokorb (2016)). Let $N$ be a Poisson random measure on the space $(0,\infty]\times S_{n-1}^{+}$, where $S_{n-1}^{+}$ is the intersection of the unit sphere and the positive orthant.
Let $N$ be a Poisson random measure on $(0,\infty]\times S_{n-1}^{+}$ with intensity measure $\zeta(\cdot)=\nu_{\alpha}(\cdot)\times{\bf P}(\Theta\in\cdot)$, where $\nu_{\alpha}$ has density $\alpha u^{-\alpha-1}$. This should be interpreted as follows: if $F$ is a subset of $(0,\infty]\times S_{n-1}^{+}$, then $N(F)$ has a Poisson distribution with parameter $\zeta(F)$.
Under our assumptions, the point process (with $\epsilon$ being the Dirac measure)
@@
is converging in distribution to $N$; this is a direct application of Theorem 6.2 of Resnick (2007). Similar to (${\rm{SCP}\textnormal{-}\rm{LT}}_{k}$) we perform a change of variables $x=y/\bar{F}^{-1}(1/k)$, exploit the homogeneity of $\phi$ and define the random variable $V_{SCP}(k)$ to be the optimal value of the optimization problem
@@
$V_{SCP}(k)$ converges weakly to the optimal value $V^{*}$ of the optimization problem
@@ 5. Example for light-tailed distributions
While the heavy-tail case directly leads to the convex program ($\rm{HT}$), the light-tail case leads to a bi-level optimization problem involving the function $I$ in (9), making it less straightforward to parse. In this section, we provide a computationally tractable example for the light-tail case that also provides insights into how the tail behavior of $L$ and its dependence structure may impact the optimal solution of ($\rm{CCP}_{\delta}$) as $\delta\downarrow 0$. Throughout this section, we assume $d=1,m=n$ and consider the problem
@@
This behavior is even worse compared to $\beta<1$ as $\delta\downarrow 0$.
@@ 6. Concluding comments
Under the light-tail assumption, the main takeaway from Section 3 is that the CVaR formulation is a good approach to chance-constrained optimization in a rare-event regime. The scenario approach and the large-deviations approximation (both suitably modified as explained in Section 3) can be a good approach to CC optimization, though for those methods, there is a nontrivial trade-off between feasibility and optimality.
@@
There are several natural research questions that may be considered: (i) we expect that large-deviations approximations, for both light-tails and heavy-tails may still yield convex programming problems for functions $\phi$ which may not be bilinear (ii) computational considerations, it will be of interest to consider scaling regimes where also the dimension of the decision variable $x$ and/or the dimension of $L$ may grow; we kept it fixed in this study; (iii) the large-deviations insights we developed could be used to develop new algorithms for chance-constrained optimization which are asymptotically optimal in the rare event regime, and also produce close to optimal solutions for fixed values of $\delta$, just like the notion of asymptotic efficiency has led to novel algorithms in the area of rare-event simulation.
@@ 7. Proofs
We prove our theorems using the following generic convergence result. There are related results in the literature, cf.  Bonnans and Shapiro (2000), Rockafellar and Wets (1998), but our direct approach seems the most efficient for our purposes.

### v2

@@ 4. Heavy tails
Throughout this section, we suppose that $L$ has a distribution that is of multivariate regular variation, which covers the case in which the elements of $L$ have a power law. To describe this assumption in a convenient form, we use a representation which (apart from not taking logarithms) is different from the one in Section 3, building on an equivalent characterization of regular variation developed in Basrak et al. (2002)). Recall that $|\cdot|$ denotes the $\ell_{1}$ norm. Our characterization guarantees that for some $\alpha>1$, and some slowly-varying function $h$,
$\bar{F}(r)={\bf P}(|L|>r)=h(r)r^{-\alpha}.$
We assume that $\bar{F}(t)$ is continuous and strictly decreasing (it has no atoms) to reduce technicalities. As a result, the inverse $\bar{F}^{-1}$ is well-defined. Moreover, our assumption on multivariate regular variation implies that for some random vector $\Theta$ defined on the $n$-dimensional unit sphere
@@
The idea behind this representation is that, if we transform $L$ in polar coordinates, the radius and angle become independent, conditional upon the radius getting large. This formulation allows dependence between extreme values of risk factors. For background on multivariate regular variation we refer to Basrak et al. (2002), Resnick (2007). While this setting includes the important example of Pareto tails, the case of log-normal tails is not included. The case of heavy-tailed Weibull tails ($\exp(-r^{\alpha}),\alpha\in(0,1)$) is covered by the framework of Section 3.
As in the light-tailed case, we first construct an appropriately normalized version of ($\rm{CCP}_{\delta}$) in the regime $\delta\downarrow 0$. We make the change of variables $x=y/r_{\delta}$, with $r_{\delta}:=\bar{F}^{-1}(\delta)=\inf\{r:{\bf P}(|L|>r)\leq\delta\}$, and write the constraint in ($\rm{CCP}_{\delta}$) as ${\bf P}(\phi(y,L)>r_{\delta})\leq\delta.$ Our assumptions that $\bar{F}(t)$ is continuous and decreasing make the chance constraint equivalent to
@@
$\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }{\bf E}[\phi(y,\Theta)^{\alpha}]\leq 1.$ ($\rm{HT}$)
@@
If $y^{*}$ solves ($\rm{HT}$), then $y^{*}/r_{\delta}$ is asymptotically optimal and $v_{\delta}\sim c^{T}y^{*}/r_{\delta}$ as $\delta\downarrow 0$.
Unlike in the light-tail case, where the optimal value of ($\rm{LT}$) could be asymptotically infeasible, the value $y^{*}$ leads to an asymptotically optimal solution as asymptotic feasibility is guaranteed.
We proceed to the approximations of ($\rm{CVaR}_{\delta}$). In contrast to the light-tailed case, in the heavy-tailed case it is no longer asymptotically optimal:
@@
$v_{{\bf CVaR},\delta}\sim v_{\delta}(1-1/\alpha)$ as $\delta\downarrow 0$.
The reason behind the subobtimality is that the overshoot $\phi(y,L)-r_{\delta}$, conditioned on $\phi(y,L)>r_{\delta}$, is not of smaller order than $r_{\delta}$ when $\phi(y,L)$ has a heavy tail. Instead, it is of the order $r_{\delta}\frac{\alpha}{\alpha-1}$.
We proceed to the scenario-based approach. In the previous section on the light-tailed case, we observed that the approximation ($\mathrm{SCP}_{k}$) of ($\rm{CCP}_{\delta}$) has some optimality properties. These properties are inherited by the convergence of the space of feasible solutions, after appropriate normalization, to a deterministic set which coincides with the feasible region of the limiting optimization problem ($\rm{LT}$). In the heavy-tail case, this is no longer true. In contrast, a random solution emerges, making any optimality and feasibility property elusive.
To see this, consider a simple example, where the goal is to maximize a scalar $x$ subject to the constraint that ${\bf}P(xL>1)\leq\delta$, with $L$ also scalar. ($\mathrm{SCP}_{k}$) can now be formulated as maximizing $x$ subject to $x\max_{i=1}^{k}L^{(i)}\leq 1$. If $L$ is light-tailed, we can normalize $\max_{i=1}^{k}L^{(i)}$ such that it converges to a deterministic constant (we have done this in a more general setting to obtain Theorem 3). However, in the heavy-tail case, such a normalization is not possible. In fact, $\max_{i=1}^{k}L^{(i)}/\bar{F}^{-1}(1/k)$ converges in distribution to a Fréchet-distributed random variable, see Resnick (2007), Nair et al. (2022) for background.
This scalar case provides the main insights into the sample approximation method in the heavy-tailed case. For completeness, we describe a general technical result using point-process theory commonly used in extreme-value theory, cf. Resnick (2007), as well as random set theory as developed in (Molchanov and Strokorb (2016)). Let $S_{n-1}^{+}$ be the intersection of the unit sphere and the positive orthant. Let $N$ be a Poisson random measure on $(0,\infty]\times S_{n-1}^{+}$ with intensity measure $\zeta(\cdot)=\nu_{\alpha}(\cdot)\times{\bf P}(\Theta\in\cdot)$, where $\nu_{\alpha}$ has density $\alpha u^{-\alpha-1}$. This should be interpreted as follows: if $F$ is a subset of $(0,\infty]\times S_{n-1}^{+}$, then $N(F)$ has a Poisson distribution with parameter $\zeta(F)$. Note that $N$ has infinitely many points, but for every $\varepsilon>0$, the number of points in $N\big((\varepsilon,\infty]\times S_{n-1}^{+}\big)$ is a.s. finite.
Under the heavy-tailed assumptions on $L$, the point process (with $\epsilon$ being the Dirac measure)
@@
is converging in distribution to $N$; this is a direct application of Theorem 6.2 of Resnick (2007).
In Section 8.6, we apply this convergence result to analyze the sample approximation method in the heavy-tailed case. Similar to (${\rm{SCP}\textnormal{-}\rm{LT}}_{k}$) we perform a change of variables $x=y/\bar{F}^{-1}(1/k)$, exploit the homogeneity of $\phi$ and define the random variable $V_{SCP}(k)$ to be the optimal value of the optimization problem
@@
$V_{SCP}(k)$ converges in distribution to the optimal value $V^{*}$ of the problem
@@
Recall that $N$ is a random point measure. The constraints $\phi(y,z)\leq 1$ in (15) are required to hold for all random points $z$ in $N$; as such, there are infinitely many constraints. However, there are only finitely many points $z$ with absolute value $\varepsilon$ or larger. Thus, it is in principle possible to verify where a specific point $y$ is feasible or not: due to the homogeneity of $\phi$ one can take this $\varepsilon$ proportional to $1/|y|$. The value $V^{*}$ of the optimization problem is random in general, as shown above in the simplest case where $y$ and $L$ are both scalar.
@@ 5. Example for light-tailed distributions
While the heavy-tail case directly leads to the convex program ($\rm{HT}$), the light-tail case leads to a bi-level optimization problem involving the function $I$ in (9), making it less straightforward to parse. The example here gives insights into how the tail behavior of $L$ and its dependence structure may impact the optimal solution of ($\rm{CCP}_{\delta}$) as $\delta\downarrow 0$.
Throughout this section, we assume $d=1,m=n$ and consider the problem
@@
This behavior is even worse compared to $\beta<1$ as $\delta\downarrow 0$. Thus, to ensure a solution is feasible for large $r$, one may take $(1-\gamma)y^{*}$ for a small $\gamma$, resulting in a solution that is close ($\gamma c^{T}y$) to the optimal value.
6. Example and Numerical Experiments
To complement our theoretical findings, we now perform numerical experiments for light- and heavy-tailed settings, demonstrating how the asymptotic predictions manifest in a practical example.
Formulation. We consider an investment portfolio optimization problem in which an investor is interested in shorting stocks that are considered overvalued. There are several websites that can be accessed to list the top 10 overvalued stocks in various industries (e.g. see tradingview, seekingalpha, etc). We do not promote any particular set of stocks to short-sell, and mention readily available sources to stress the applicability of this problem formulation.
We consider an investment horizon, for example, one quarter of a year. The quantity $x_{i}$ corresponds to the number of shares that we will short-sell of the $i$-th asset/stock. As it is commonly done in practice, we relax integer constraints. Let $s_{i}$ be the price of a single share of the $i$-th stock today and $S_{i}$ be the value of the $i$-th stock’s share in a quarter of a year. Prices are expressed in units of monetary risk relative to the Value-at-Risk that the manager is willing to tolerate within a given confidence level. The Value-at-Risk at level $1-\delta$ is the largest amount that can be lost with probability at most $1-\delta$. For instance, if the portfolio manager is willing to tolerate at most $\mathdollar 10K$ of risk (measured by the Value-at-Risk at 95% level) and a single share of the first asset is worth $\mathdollar 1K$, then $s_{1}=0.1$ and $\delta=0.05$. Since we are assuming that the stock is overvalued, $\mu_{i}={\bf E}(S_{i})-s_{i}<0$ and we write $c_{i}=-\mu_{i}$. Define $L_{i}=S_{i}-s_{i}$ and note that the expected return earned by shorting the stocks according to the allocation $x=(x_{1},...,x_{n})^{T}$ at the end of a quarter of a year is precisely $c^{T}x$ (which is to be maximized). The amount of money that we can lose, on the other hand, is $x^{T}L$. So, we are interested in maximizing the expected return subject to the Value-at-Risk at level $1-\delta$ is not larger than 1. We may add a budget constraint, say $x^{T}s\leq b_{0}$, but we will ignore this, as this constraint is not binding as $\delta\downarrow 0$. Precisely, we are interested in solving
$\mbox{maximize }c^{T}x\hskip 28.45274pt\mbox{ subject to }x\geq 0,\mbox{ and }{\bf P}\bigg(\sum_{i=1}^{n}x_{i}L_{i}\geq 1\bigg)\leq\delta,$
In order to test the quality of our approximations, we will use a class of models for $L$ for which this problem can be efficiently solved. We consider two types of models for $L$, one light-tailed and another one is heavy-tailed, so we can test the performance of the corresponding approximations we developed for each of these two cases. We choose $n=10$ assets. For the light-tailed case, $L$ is assumed to be Gaussian, and for the heavy-tailed case $L$ is $t$-distributed with $\alpha=5$ degrees of freedom. In both cases, $L$ has the same mean and covariance parameters, which have been selected at random; the means are uniformly distributed in [-10,1], the covariance matrix was sampled from a Wishart model, with mean identity and 13 degrees of freedom (this ensures a positive definite covariance matrix). The choice of $k$ for the scenario approach is
$k(\delta,0.05)=\left\lceil\frac{2n\log(2/\delta)}{\delta}+\frac{2\log(1/.05)}{\delta}+2n\right\rceil.$
Results. Our theoretical results are derived for non-negative $L$, which is a technical assumption imposed to facilitate our development. While we choose our models for tractability, our choice also stresses that our insights apply more broadly [^1] The experiments were performed in Matlab 2024, for reproducibility, the random seed was selected as rng(42)..
The results are shown in Figure 1, organized in two rows and two columns. The first row corresponds to the light-tailed case and the second row corresponds to the heavy-tailed case. The first column corresponds to what we call “moderate confidence” chance constraint (enforcing values of $\delta\in[10^{-3},5\times 10^{-5}]$) and the second column corresponds to what we call “high confidence” chance constraint (enforcing values of $\delta\in[10^{-4},10^{-7}]$). The intervals have been chosen overlapping so we can appreciate continuity in the behavior.
Each of the four figures in the panel has two vertical axis, the left vertical axis shows the estimated optimal expected return (i.e. the optimal value of the optimization problem) under the different methods. The actual value of the chance-constraint optimization problem is denoted as VaR. We also include the CVaR approximation, the scenario approach, and the large deviations approximation. The right vertical axis plots the ratios of the CVaR to the VaR solutions. In the light-tailed case this ratio converges to 1 according to our theory, whereas in the heavy-tailed case the ratio converges to the ratio $(1-1/\alpha)=.8$ (in our case). These limiting theoretical values are represented as the “Limit”. The plots show that convergence eventually occurs as the required confidence increases (i.e. the probability of chance constrain violation decreases).
In the high confidence case, in both cases, the worst performance is obtained by the scenario approach, the large deviations approximations dominates the scenario approach. The performance of the large deviations approximation is close to that of the scenario approach in the light tailed case, but significantly better in the heavy tailed case; the CVaR relaxation is the best performing approximation among all of the options that we consider.
In the very high confidence case, we do not present the scenario approach because it was too time consuming to produce. We can observe, as predicted by the theory, that the large deviations approximation overtakes the CVaR approximation in terms of quality of performance. The ratio VaR/CVaR is close to the predicted limit, especially for the heavy-tailed case.
The scenario approach is considered a viable and highly popular alternative, the large deviations approximations exhibits comparable performance in the light-tailed case but performs significantly better in the heavy-tailed case, with a significantly lower computational overhead. The CVaR is a viable relaxation, but eventually (in high reliability regimes) is overtaken by the large deviations approximations.
(a) Light Tails - Moderate Confidence
(b) Light Tails - High Confidence
(c) Heavy Tails - Moderate Confidence
(d) Heavy Tails - High Confidence
Figure 1. A detailed explanation of the figures is given when Figure 1 is first introduced. The left vertical axis shows the exact chance constrained solution (represented as VaR) compared to the CVaR relaxation, the scenario approach and the large deviations approximation; the right vertical axis shows the ratios of CVaR/VaR solutions which converge to the theoretical limit (dotted line) as $\delta\downarrow 0$. The figures on the right can be seen as a continuation of the figures on the left as the values of $\delta$ decrease.
@@ 7. Concluding comments
Supported by the numerical experiments, the theoretical results in this paper provide a consistent picture of how chance-constrained programs behave in the rare-event regime. We summarize our findings before proceeding to the proofs.
@@
Under the light-tail assumption, the main takeaway from Section 3 is that the CVaR and scenario formulations —with carefully chosen number of scenarios— are good approaches to chance-constrained optimization in a rare-event regime. Their solutions are asymptotically feasible (with probability tending to one), and are asymptotically optimal. The large-deviations approximation (suitably modified as explained in Section 3) can be a good approach to chance-constrained optimization and leads to an asymptotically accurate estimate of the optimal value, though feasibility of its solution cannot be guaranteed.
@@
There are several natural research questions that may be considered: (i) we expect that large-deviations approximations, for both light-tails and heavy-tails may still yield convex programming problems for functions $\phi$ which are not be bilinear; (ii) computationally, it will be of interest to consider scaling regimes where also the dimension of the decision variable $x$ and/or the dimension of $L$ may grow —it is kept fixed in this study; (iii) the large-deviations insights we developed could be used to develop new algorithms for chance-constrained optimization which are asymptotically optimal in the rare event regime, and also produce close to optimal solutions for fixed values of $\delta$, just like the notion of asymptotic efficiency has led to novel algorithms in the area of rare-event simulation.
@@ 8. Proofs
We first prove Proposition 1. Afterwards, we state and prove a preliminary that we use in most of our theorems that we prove afterwards.
Proof of Proposition 1
We first show convexity. Let $b_{1},b_{2}$ be such that $I(b_{j})\geq 1,j=1,2$. Let $a\in(0,1)$. We have to show that $I(ab_{1}+(1-a)b_{2})\geq 1$. Since $\lambda$, defined in (7), is continuous, non-decreasing, and satisfies the relation $\lambda(rx)=r^{\beta}\lambda(x)$, there exists an $x_{a}\geq 0$ such that $I(ab_{1}+(1-a)b_{2})=\lambda(x^{(a)})$. Moreover, by definition of $I$ in (9), there exists a sequence $(y_{\varepsilon})_{\varepsilon>0}$ such that $(ab_{1}+(1-a)b_{2})^{T}y_{\varepsilon}\geq 1$ and $\lambda(y_{\varepsilon})\downarrow\lambda(x^{(a)})$ as $\varepsilon\downarrow 0$. For each $\varepsilon>0$, $\max_{j=1,2}b_{j}^{T}y_{\varepsilon}\geq 1$. Suppose without loss of generality that $b_{j}^{T}x^{(a)}\geq 1$. Then, for some $\delta_{\varepsilon}$ tending to 0 as $\varepsilon\downarrow 0$,
$I(ab_{1}+(1-a)b_{2})\geq\lambda(x^{(a})-\delta_{\varepsilon}\geq\inf_{x:b_{1}^{T}x\geq 1}\lambda(x)-\delta_{\varepsilon}=I(b^{(1)})-\delta_{\varepsilon}\geq 1-\delta_{\varepsilon}.$
Since this is true for each $\varepsilon>0$, $I(ab^{(1)}+(1-a)b^{(2)})\geq 1$, which implies the desired convexity property.
Next, we show that $I(\cdot)$ is continuous, which implies that $\{b:I(b)\geq 1\}$ is closed. We apply Theorem 1.17(c) of Rockafellar and Wets (1998). Following their notation, we write $f(x,u)=\lambda(x)$ if $u^{T}x\geq 1$ and $f(x,u)=\infty$ otherwise. This function is lower-semicontinuous. In addition, given $\bar{u}\geq 0$ and $M\in[0,\infty)$, there exists a neighborhood $V$ of $\bar{u}$ such that $\{(x,u):u\in V,f(x,u)\leq M\}$ is bounded, since $\lambda(x)=|x|^{\beta}\lambda(x/|x|)$ and $\inf_{x:|x|=1}\lambda(x)>0$ by (8). Thus, $f$ is uniformly level bounded conform Definition 1.16 in Rockafellar and Wets (1998). Let $p(u):=\inf_{x}f(x,u)=\inf\{\lambda(x):u^{T}x\geq 1,x\in[0,\infty)^{m}\}=I(u)$. By continuity of $\lambda$, the infimum is attained at a (possibly non-unique) value $\bar{x}$ for a given $u$. Since the definition of $\bar{x}$ implies that $u^{T}\bar{x}\geq 1$, we see that $f$ is continuous at $(u,\bar{x})$, so that all conditions of Theorem 1.17(c) of Rockafellar and Wets (1998) hold. We conclude that $p(u)=I(u)$ is continuous at $u\geq 0$. Therefore $\{b:I(b)\geq 1\}$ is closed. Boundedness of $\{b:I(b)\geq 1\}$ follows from the relation $\lambda(rx)=r^{\beta}\lambda(x)$, $\beta>0$, and thus the set is compact.
We now turn to a proof of (10). We first construct a lower bound. Given $\varepsilon>0$, let $x$ be such that $b^{T}x\geq 1$ and $\lambda(x)\leq I(b)+\varepsilon$. Observe that if $L>rx$, then $b^{T}L>rb^{T}x\geq 1$ for any $b\geq 0$, which implies $\{b^{T}L>r\}\supseteq\{L>rx\}$. Using Assumption (7), it follows that, as $r\to\infty$
${\bf P}(b^{T}L>r)\geq{\bf P}(L>rx)=e^{-q(r)\lambda(x)(1+o(1))}\geq e^{-q(r)(I(b)+\varepsilon)(1+o(1))}.$ (27)
Since this holds for every $\varepsilon>0$, the lower bound of (10) follows.
We turn to the upper bound. If $b$ contains elements with zeroes, we may leave out the corresponding coordinates of $L$, and the reduced random vector $\tilde{L}$ still satisfies Assumption (7), with function $\lambda((0,x))$ (with zeroes at the corresponding coordinates of $b$). So, we may assume without loss of generality that the coordinates of $b$ are all strictly positive. As a result, given $\varepsilon>0$, the set $A_{\varepsilon}=\{x\geq 0:b^{T}x=1-\varepsilon\}$ is compact. Let $\{x_{1},\ldots,x_{N}\}\subset A_{\varepsilon}$, for some large constant $N=N(\varepsilon)$, be constructed such that
$\{y\geq 0:b^{T}y>1\}\,\subseteq\,\bigcup_{j\in[N]}\{y>x_{j}:b^{T}y>1\}\,\,\,\Big(\subseteq\,\big\{y>x_{j}\text{ for some }j\in[N]\big\}\Big),$
which is possible when $N$ is sufficiently large. We obtain by a union bound and (7), as $r\to\infty$,
${\bf P}\big(b^{T}(L/r)>1)\leq\sum_{j=1}^{N}{\bf P}(L/r>x_{j})\leq N\max_{j}{\bf P}(L>rx_{j})$ $=Ne^{-q(r)(1+o(1))\min_{j}\lambda(x_{j})}.$
Since all $x_{j}$ satisfy $b^{T}x_{j}=1-\varepsilon$, we have $\lambda(x_{j})\geq I(b/(1-\varepsilon))=(1-\varepsilon)^{\beta}I(b)$. Thus, as $r\to\infty$,
${\bf P}(b^{T}L>r)\leq e^{-q(r)(1+o(1))(1-\varepsilon)^{\beta}I(b)},$
providing a matching upper bound to (27) and thus proving (10). Lastly, we use the principle of the largest term (see e.g. Ganesh et al. (2004), Lemma 2.1) to conclude the asymptotics in (11). $\square$
We proceed to a preliminary result that is used in the convergence results of Theorems 134, and 5. There are related results in the literature, cf. Bonnans and Shapiro (2000), Rockafellar and Wets (1998), but our direct approach seems the most efficient for our purposes. For $x,y\in\mathbb{R}^{m}$, we say that $x\leq y$ if $x_{i}\leq y_{i}$ for all $i\in[m]$. If one of these inequalities is strict, we write $x<y$. We call $f:\mathbb{R}^{m}\mapsto\mathbb{R}$ (strictly) increasing if $x\leq y$ (resp. $x<y$) implies that $f(x)\leq f(y)$ (resp. $f(x)<f(y)$).